OpenStudy (anonymous):

help?!? will give a medal! Find the area of a circle with a diameter of 22 inches. Use 3.14 for pi. 379.94 in2 452.16 in2 527.52 in2 621.72 in2

OpenStudy (anonymous):

do you know the equation of a circle?

OpenStudy (anonymous):

sorry, equation for the area of a circle

OpenStudy (anonymous):

are you there luka?

OpenStudy (anonymous):

@esshotwired

OpenStudy (esshotwired):

do you know the equation for the area of a circle?

OpenStudy (anonymous):

A = π × r2 A = (π/4) × D2

Directrix (directrix):

A = π × r2 A = 3.14 * (11)² --> if the diameter is 22, radius is 22/2 A = ? @luka1998 Post what you get, okay?

OpenStudy (anonymous):

379.94

Directrix (directrix):

That is what I got, too.

OpenStudy (anonymous):

thx @Directrix

Directrix (directrix):

Always a pleasure to work with you, @luka1998 You are welcome.

OpenStudy (anonymous):

can u help me in another question ?

OpenStudy (anonymous):

The volume of a fish tank is 40 cubic feet. If the density is 0.4fish over feet cubed, how many fish are in the tank? 16 4 100 10